### v2

# Classifying localizing subcategories of a Grothendieck category

Abstract.
Let $\mathcal{A}$ be a locally coherent Grothendieck category, $\operatorname{fp-}\mathcal{A}$ be the full subcategory of $\mathcal{A}$ consisting of finitely presented objects and $\operatorname{ASpec}\mathcal{A}$ be the atom spectrum of $\mathcal{A}$. In this paper, we classify localizing subcategories of finite type of $\mathcal{A}$ via open subsets of $\operatorname{ASpec}\mathcal{A}$. We investigate $\operatorname{ASpec}\operatorname{fp-}\mathcal{A}$ and show that if $\operatorname{ASpec}\operatorname{fp-}\mathcal{A}=\operatorname{ASpec}\mathcal{A}$, then $\mathcal{A}$ is locally noetherian. As an application, we specialize our investigation to the case of commutative coherent rings.
Key words and phrases:
Atom spectrum; coherent ring, localizing subcategory; Serre subcategoy; Ziegler spectrum.
2020 Mathematics Subject Classification:
13C05, 18E10, 18E35
1. Introduction
Throughout this paper, we assume that $\mathcal{A}$ is a Grothendieck category. The idea of classifying subcategories of $\mathcal{A}$ using an appropriate spectrum originates from the classification of subcategories of modules over a commutative ring $A$ using $\operatorname{Spec}A$, which has been established by numerous authors (e.g. [Hop, Hov, GP, T]). The initial work in extending this concept to Grothendieck categories was carried out by Gabriel [G], who defined the spectrum $\operatorname{Sp}\mathcal{A}$ as the set of isomorphism classes of indecomposable injective objects in $\mathcal{A}$. In the case of a commutative noetherian ring, these indecomposable injective modules are in one-to-one correspondence with prime ideals.
For a locally coherent category $\mathcal{A}$, Herzog [H] defined $\operatorname{Zg}\mathcal{A}$, the Ziegler spectrum of $\mathcal{A}$, which imposes an alternative topology on $\operatorname{Sp}\mathcal{A}$. This topology was originally defined by Ziegler [Z] associating to a ring $A$, a topological space whose points are the isomorphism classes of indecomposable pure-injective $A$-modules.
Recently, Kanda [K] introduced the atom spectrum $\operatorname{ASpec}\mathcal{A}$, for an abelian category $\mathcal{A}$ that may not have enough injective objects, and endowed it with a topology. This construction was inspired by Storrer’s [St] work on monoform modules and their equivalence relation over non-commutative rings. For an object $M$ in $\mathcal{A}$, the atom support $\operatorname{ASupp}M$ can be defined analogously to the support of a module over a commutative ring. This paper is devoted to classifying the localizing subcategories of finite type of a locally coherent category $\mathcal{A}$ via the open subsets of $\operatorname{ASpec}\mathcal{A}$.
Unfortunately, Grothendieck categories does not have generally enough atoms, which limits our ability to find out further insights about $\mathcal{A}$ via $\operatorname{ASpec}\mathcal{A}$. A well-known example of a categories with enough atoms is locally noetherian categories; see [K]. An abelian category $\mathcal{A}$ is said to be locally monoform if every nonzero object in $\mathcal{A}$ contains a monoform subobject. In Section 2, we show that if $\mathcal{A}$ is a locally monoform category, then there is a bijective correspondence between the lattice of localizing subcategories of $\mathcal{A}$ and the lattice of open subsets of $\operatorname{ASpec}\mathcal{A}$; see section˜2 . Furthermore, we also show that semi-noetherian categories are locally monoform; see section˜2.
A Grothendieck category $\mathcal{A}$ is said to be locally coherent if $\mathcal{A}$ has a generating set of finitely presented objects and the full subcategory $\operatorname{fp-}\mathcal{A}$ of finitely presented objects in $\mathcal{A}$ is abelian. In Section 3, we assume that $\mathcal{A}$ is locally coherent. We define a new topology on $\operatorname{ASpec}\mathcal{A}$ in which $\{\operatorname{ASupp}M|\hskip 2.84544ptM\in\operatorname{fp-}\mathcal{A}\}$ forms a basis of open subsets of $\operatorname{ASpec}\mathcal{A}$. We use the symbol $\operatorname{ZASpec}\mathcal{A}$ instead of $\operatorname{ASpec}\mathcal{A}$ with this topology. For any subcategory $\mathcal{X}$ of $\mathcal{A}$ and any subset $\mathcal{U}$ of $\operatorname{ASpec}\mathcal{A}$, assume that $\operatorname{ASupp}\mathcal{X}=\bigcup_{M\in\mathcal{X}}\operatorname{ASupp}M$ and $\operatorname{ASupp}^{-1}\mathcal{U}=\{M\in\mathcal{A}|\hskip 2.84544pt\operatorname{ASupp}M\subseteq\mathcal{U}\}$. The following theorem classifies Serre subcategories of $\operatorname{fp-}\mathcal{A}$ using open subsets of $\operatorname{ZASpec}\mathcal{A}$:
Theorem 1.1 (section˜3).
Let $\mathcal{A}$ be a locally coherent category. Then the map $\mathcal{X}\mapsto\operatorname{ASupp}\mathcal{X}$ provides a bijective correspondence between the lattice of Serre subcategories of $\operatorname{fp-}\mathcal{A}$ and the lattice of open subsets of $\operatorname{ZASpec}\mathcal{A}$. The inverse map is given by $\mathcal{U}\mapsto\operatorname{ASupp}^{-1}\mathcal{U}\cap\operatorname{fp-}\mathcal{A}$.
As a conclusion, we prove that there is a bijective correspondence between the lattice of open subsets of $\operatorname{ZASpec}\mathcal{A}$ and the lattice of localizing subcategories of finite type of $\mathcal{A}$; see section˜3. Furthermore, we show that a localizing subcategory $\mathcal{X}$ is of finite type of $\mathcal{A}$ if and only if $\operatorname{ASupp}\mathcal{X}$ is an open subset of $\operatorname{ZASpec}\mathcal{A}$. As $\operatorname{fp-}\mathcal{A}$ is abelian, $\operatorname{ASpec}\operatorname{fp-}\mathcal{A}$ can be investigated independently. It is known as Cohen’s theorem that if prime ideals of a commutative ring $A$ are finitely generated, then $A$ is noetherian. In the following theorem we extend this theorem for locally coherent categories:
Theorem 1.2 (section˜3).
Let $\mathcal{A}$ be a locally coherent category. If $\operatorname{ASpec}\operatorname{fp-}\mathcal{A}=\operatorname{ASpec}\mathcal{A}$, then $\mathcal{A}$ is locally noetherian.
We denote the atoms in $\operatorname{ASpec}\operatorname{fp-}\mathcal{A}$ by $\alpha_{f}=\overline{H}$, where $H$ is a monoform object in $\operatorname{fp-}\mathcal{A}$. It is shown that $H$ is also a monoform object in $\mathcal{A}$ so that $\alpha=\overline{H}$ is an atom in $\operatorname{ASpec}\mathcal{A}$. Then the map $\alpha_{f}\stackrel{{\scriptstyle g}}{{\mapsto}}\alpha$ embeds $\operatorname{ASpec}\operatorname{fp-}\mathcal{A}$ as a subspace of $\operatorname{ASpec}\mathcal{A}$. Then for any object $M$ in $\mathcal{A}$, we define $\operatorname{fASupp}M=\{\alpha_{f}=\overline{H}\in\operatorname{ASpec}\operatorname{fp-}\mathcal{A}|\hskip 2.84544pt\alpha=\overline{H}\in\operatorname{ASupp}M\}.$ For any subcategory $\mathcal{X}$ of $\mathcal{A}$, we define $\operatorname{fASupp}\mathcal{X}=\bigcup_{M\in\mathcal{X}}\operatorname{fASupp}M$. In the following theorem, we provide a sufficient condition under which a localizing subcategory of $\mathcal{A}$ is of finite type:
Theorem 1.3 (section˜3).
Let $\mathcal{A}$ be a locally coherent category. Then any localizing subcategory $\mathcal{X}$ of $\mathcal{A}$ with $\operatorname{fASupp}\mathcal{X}=\operatorname{ASupp}\mathcal{X}$ is of finite type.
For any object $M$ in $\mathcal{A}$, set $\operatorname{\mathcal{O}}(M)=\{I\in\operatorname{Zg}\mathcal{A}|\hskip 2.84544pt\operatorname{Hom}(M,I)\neq 0\}.$ Herzog showed that over a locally coherent category, the set $\{\operatorname{\mathcal{O}}(M)|\hskip 2.84544ptM\in\operatorname{fp-}\mathcal{A}\}$ can serve as a basis for a topology on $\mathcal{A}$. This topology, denoted by $\operatorname{Zg}\mathcal{A}$, is known as Ziegler topology of $\mathcal{A}$. A category $\mathcal{A}$ is said to be locally uniform if every nonzero object in $\mathcal{A}$ has a uniform subobject. In Section 4, we show that $\mathcal{A}$ is locally monoform if and only if $\mathcal{A}$ is locally uniform and the map $\theta:\operatorname{ZASpec}\mathcal{A}\rightarrow\operatorname{Zg}\mathcal{A}$, given by $\alpha\mapsto E(\alpha)$ is a homeomorphism. We provide an example which show that any locally coherent is not necessarily locally monoform; see section˜4. Furthermore, we provide a sufficient condition under which a locally coherent category is semi-noetherian (and so a locally monoform); see section˜4.
In Section 5, we assume that $A$ is a commutative coherent ring. We denote $\operatorname{ZASpec}\operatorname{Mod}A$ and $\operatorname{Zg}\operatorname{Mod}A$ by $\operatorname{ASpec}A$ and $\operatorname{Zg}A$. For a spectral topological space $X$, Hochster [Ho] endows the underlying set with a new, dual topological by defining its open subsets as those of the form $Y=\bigcup_{i\in\varOmega}Y_{i}$, where $X\setminus Y_{i}$ is a quasi-compact open subset of $X$ for each $i\in\varOmega$. The symbol $X^{\star}$ denotes $X$ with the new topology. We write $\operatorname{Spec}^{\star}A$ for $(\operatorname{Spec}A)^{\star}$. We establish a bijective correspondence between the lattice of open subsets of $\operatorname{ZASpec}^{\star}A$ and that of $\operatorname{ZASpec}A$; see section˜5. As our main application, we use Gabriel topologies on $A$ to classify the localizing subcategories of modules of finite type over a commutative coherent ring:
Theorem 1.4 (section˜5).
There are bijective correspondence between:
$\rm(1)$} the lattice of open subsets of $\operatorname{ZASpec}A$,
$\rm(2)$} the lattice of closed subsets of $\operatorname{Zg}A$,
$\rm(3)$} the lattice of Gabriel topologies having bases of finitely generated ideals of $A$,
$\rm(4)$} The lattice of localizing subcategory of finite type of $\operatorname{Mod}A$.
2. Categories with enough atoms
We begin this section with key definitions which are used throughout this paper.
Definition 2.1.
(1) An abelian category $\mathcal{A}$ with a generator is said to be a Grothendieck category if it has arbitrary direct sums and direct limits of short exact sequence are exact, this means that if a direct system of short exact sequences in $\mathcal{A}$ is given, then the induced sequence of direct limits is a short exact sequence.
(2) An object $M$ in $\mathcal{A}$ is finitely generated if whenever there are subobjects $M_{i}\leq M$ for $i\in I$ satisfying $M=\underset{i\in I}{\varSigma}M_{i}$, then there is a finite subset $J\subseteq I$ such that $M=\underset{i\in J}{\varSigma}M_{i}$. A category $\mathcal{A}$ is said to be locally finitely generated if it has a small generating set of finitely generated objects.
(3) A category $\mathcal{A}$ is said to be locally noetherian if it has a small generating set of noetherian objects.
Definition 2.2.
(i) An object $U$ in $\mathcal{A}$ is said to be uniform if every nonzero subobject of $U$ is an essential subobject of $U$.
(ii) A nonzero object $M$ in $\mathcal{A}$ is monoform if for any nonzero subobject $N$ of $M$, there exists no common nonzero subobject of $M$ and $M/N$ which means that there does not exist a nonzero subobject of $M$ which is isomorphic to a subobject of $M/N$. We denote by $\operatorname{ASpec}_{0}\mathcal{A}$, the set of all monoform objects in $\mathcal{A}$.
Two monoform objects $H$ and $H^{\prime}$ in $\mathcal{A}$ are said to be atom-equivalent if they have a common nonzero subobject. The atom equivalence establishes an equivalence relation on monoform objects. For every monoform object $H$ in $\mathcal{A}$, we denoted by $\overline{H}$, the equivalence class of $H$, that is
$\overline{H}=\{G\in\operatorname{ASpec}_{0}\mathcal{A}|\hskip 2.84544ptH\hskip 2.84544pt{\rm and}\hskip 2.84544ptG$ have a common nonzero subobject}.\}.
Definition 2.3.
The atom spectrum $\operatorname{ASpec}\mathcal{A}$ of $\mathcal{A}$ is the quotient class of $\operatorname{ASpec}_{0}\mathcal{A}$ consisting of all equivalence classes induced by this equivalence relation; in other words
$\operatorname{ASpec}\mathcal{A}=\{\overline{H}|\hskip 2.84544ptH\in\operatorname{ASpec}_{0}\mathcal{A}\}.$
Any equivalence class is called an atom of $\operatorname{ASpec}\mathcal{A}$. We observe that since $\mathcal{A}$ is a Grothendieck category, $\operatorname{ASpec}\mathcal{A}$ is a set.
The notion support and associated prime of modules over a commutative ring can be generalized to objects in a Grothendieck category $\mathcal{A}$ as follows:
Definition 2.4.
Let $M$ be an object in $\mathcal{A}$.
(1) The atom support of $M$, denoted by $\operatorname{ASupp}M$, is defined as
$\operatorname{ASupp}M=\{\alpha\in\operatorname{ASpec}\mathcal{A}|\hskip 2.84544pt{\rm there\hskip 2.84544ptexists\hskip 2.84544pt}H\in\alpha\hskip 2.84544pt{\rm which\hskip 2.84544ptis\hskip 2.84544pta\hskip 2.84544ptsubquotient\hskip 2.84544ptof\hskip 2.84544pt}M\}.$
(2) The associated atom of $M$, denoted by $\operatorname{AAss}M$, is defined as
$\operatorname{AAss}M=\{\alpha\in\operatorname{ASupp}M|\hskip 2.84544pt{\rm there\hskip 2.84544ptexists\hskip 2.84544pt}H\in\alpha\hskip 2.84544pt{\rm which\hskip 2.84544ptis\hskip 2.84544pta\hskip 2.84544ptsubobject\hskip 2.84544ptof\hskip 2.84544pt}M\}.$
In view of [Sto, p. 631], for a commutative ring $A$, there is a bijection between $\operatorname{ASpec}A$ and $\operatorname{Spec}A$. A subset $\varPhi$ of $\operatorname{Spec}A$ is said to be specialization-closed if for any prime ideals $\mathfrak{p}$ and $\mathfrak{q}$ of $A$ with $\mathfrak{p}\subseteq\mathfrak{q}$, the condition $\mathfrak{p}\in\varPhi$ implies that $\mathfrak{q}\in\varPhi$. A corresponding subset in $\operatorname{ASpec}\mathcal{A}$ can be defined as follows:
Definition 2.5.
A subset $\varPhi$ of $\operatorname{ASpec}\mathcal{A}$ is said to be open if for any $\alpha\in\varPhi$, there exists a monoform object $H$ in $\mathcal{A}$ such that $\alpha=\overline{H}$ and $\operatorname{ASupp}H\subset\varPhi$. For any object $M$ in $\mathcal{A}$, it is clear that $\operatorname{ASupp}M$ is an open subset of $\operatorname{ASpec}\mathcal{A}$. Then $\operatorname{ASpec}\mathcal{A}$ is equipped with a topology having $\{\operatorname{ASupp}M|\hskip 2.84544ptM\in\mathcal{A}\}$ as a basis of open subsets. For any subcategory $\mathcal{X}$ of $\mathcal{A}$, we set $\operatorname{ASupp}\mathcal{X}=\underset{M\in\mathcal{X}}{\bigcup}\operatorname{ASupp}M$ which is an open subset of $\operatorname{ASpec}\mathcal{A}$. Also, for every subset $\mathcal{U}$ of $\operatorname{ASpec}\mathcal{A}$, we set $\operatorname{ASupp}^{-1}\mathcal{U}=\{M\in\mathcal{A}|\hskip 2.84544pt\operatorname{ASupp}M\subseteq\mathcal{U}\}$ which is a Serre subcategory of $\mathcal{A}$. Furthermore, since $\mathcal{A}$ is a Grothendieck category, it follows from [K, Proposition 3.11] that $\operatorname{ASupp}^{-1}\mathcal{U}$ is a localizing subcategory of $\mathcal{A}$.
Definition 2.6.
An atom $\alpha$ in $\operatorname{ASpec}\mathcal{A}$ is said to be maximal if there exists a simple object $H$ in $\mathcal{A}$ such that $\alpha=\overline{H}$. The class of all maximal atoms in $\operatorname{ASpec}\mathcal{A}$ is denoted by $\textnormal{m-}\operatorname{ASpec}\mathcal{A}$.
Definition 2.7.
A full subcategory $\mathcal{X}$ of an abelian category $\mathcal{A}$ is said to be Serre if for any exact sequence $0\rightarrow M\rightarrow N\rightarrow K\rightarrow 0$ in $\mathcal{A}$, the object $N$ belongs to $\mathcal{X}$ if and only if $M$ and $K$ belong to $\mathcal{X}$.
Definition 2.8.
A Serre subcategory $\mathcal{X}$ of $\mathcal{A}$ is said to be localizing if the canonical functor $F:\mathcal{A}\rightarrow\mathcal{A}/\mathcal{X}$ admits a right adjoint functor $G:\mathcal{A}/\mathcal{X}\rightarrow\mathcal{A}$ (known as the section functor). For every subcategory $\mathcal{C}$ of $\mathcal{A}$, we denote by $\left\langle\mathcal{C}\right\rangle_{\rm loc}$, the localizing subcategory of $\mathcal{A}$ generated by $\mathcal{C}$ which is the smallest localizing subcategory of $\mathcal{A}$ containing $\mathcal{C}$. The right adjoint of the inclusion functor $i:\mathcal{X}\rightarrow\mathcal{A}$, denoted $t_{\mathcal{X}}:\mathcal{A}\rightarrow\mathcal{X}$ is said to be the radical functor associated to $\mathcal{X}$. For any object $M$ in $\mathcal{A}$, $t_{\mathcal{X}}(M)$ is the largest subobject of $M$ contained in $\mathcal{X}$.
Definition 2.9.
An object $M$ in $\mathcal{A}$ is said to be locally monoform if every nonzero quotient of $M$ contains a monoform subobject. An abelian category $\mathcal{A}$ is said to be locally monoform if every nonzero object in $\mathcal{A}$ is locally monoform.
The following lemma is crucial to classify the localizing subcategories of a locally monoform category.
Lemma 2.10.
Let $\mathcal{X}$ be a localizing subcategory of $\mathcal{A}$ and let $M$ be a locally monoform object in $\mathcal{A}$. If $\operatorname{ASupp}M\subset\operatorname{ASupp}\mathcal{X}$, then $M\in\mathcal{X}$.
Proof.
Assume that $M$ is not in $\mathcal{X}$ and $t_{\mathcal{X}}(M)$ is the largest subobject of $M$ belonging to $\mathcal{X}$. Since $M$ is a locally monoform, $M/t_{\mathcal{X}}(M)$ contains a monoform subobject $N/t_{\mathcal{X}}(M)$. Therefore $\overline{N/t_{\mathcal{X}}(M)}\in\operatorname{ASupp}\mathcal{X}$ and so there exists an object $X\in\mathcal{X}$ such that $\overline{N/t_{\mathcal{X}}(M)}\in\operatorname{ASupp}X$. Thus $N/t_{\mathcal{X}}(M)$ contains a nonzero subobject isomorphic to a subquotient of $X$. But this implies that $t_{\mathcal{X}}(N/t_{\mathcal{X}}(M))$ is nonzero which is a contradiction. ∎
Proposition 2.11.
Let $\mathcal{A}$ be a locally monoform category. Then the map $\mathcal{X}\mapsto\operatorname{ASupp}\mathcal{X}$ provides a bijective correspondence between the lattice of localizing subcategories of $\mathcal{A}$ and the lattice of open subsets of $\operatorname{ASpec}\mathcal{A}$. The inverse map is given by $\mathcal{U}\mapsto\operatorname{ASupp}^{-1}\mathcal{U}$.
Proof.
For a localizing subcategory $\mathcal{X}$ of $\mathcal{A}$ and an open subset $\mathcal{U}$ of $\operatorname{ASpec}\mathcal{A}$, it suffices to prove that $\operatorname{ASupp}^{-1}(\operatorname{ASupp}\mathcal{X})=\mathcal{X}$ and $\operatorname{ASupp}(\operatorname{ASupp}^{-1}\mathcal{U})=\mathcal{U}$. But this follows easily from the definition and using section˜2. ∎
It is clear that locally noetherian categories are locally monoform. In the rest of this section we show that semi-noetherian categories are another example of locally monoform categories.
Definition 2.12.
We define the Krull-Gabriel filtration of $\mathcal{A}$ as follows:
For any ordinal (i.e ordinal number) $\sigma$ we denote by $\mathcal{A}_{\sigma}$, the localizing subcategory of $\mathcal{A}$ which is defined in the following manner:
$\mathcal{A}_{-1}$ is the zero subcategory.
$\mathcal{A}_{0}$ is the smallest localizing subcategory containing all simple objects.
Let us assume that $\sigma=\rho+1$ and denote by $F_{\rho}:\mathcal{A}\rightarrow\mathcal{A}/\mathcal{A}_{\rho}$ the canonical functor and by $G_{\rho}:\mathcal{A}/\mathcal{A}_{\rho}\rightarrow\mathcal{A}$ the right adjoint functor of $F_{\rho}$. Then an object $X$ in $\mathcal{A}$ will belong to $\mathcal{A}_{\sigma}$ if and only if $F_{\rho}(X)\in{\rm Ob}(\mathcal{A}/\mathcal{A}_{\rho})_{0}$. The left exact radical functor (torsion functor) corresponding to $\mathcal{A}_{\rho}$ is denoted by $t_{\rho}$. If $\sigma$ is a limit ordinal, then $\mathcal{A}_{\sigma}$ is the localizing subcategory generated by all localizing subcategories $\mathcal{A}_{\rho}$ with $\rho<\sigma$. It is clear that if $\sigma\leq\sigma^{\prime}$, then $\mathcal{A}_{\sigma}\subseteq\mathcal{A}_{\sigma^{\prime}}$. Moreover, since the class of all localizing subcategories of $\mathcal{A}$ is a set, there exists an ordinal $\tau$ such that $\mathcal{A}_{\sigma}=\mathcal{A}_{\tau}$ for all $\sigma\geq\tau$. Let us put $\mathcal{A}_{\tau}=\cup_{\sigma}\mathcal{A}_{\sigma}$. Then $\mathcal{A}$ is said to be semi-noetherian if $\mathcal{A}=\mathcal{A}_{\tau}$. We also say that the localizing subcategories $\{\mathcal{A}_{\sigma}\}_{\sigma}$ define the Krull-Gabriel filtration of $\mathcal{A}$. We say that an object $M$ in $\mathcal{A}$ has the Krull-Gabriel dimension defined or $M$ is semi-noetherian if $M\in{\rm Ob}(\mathcal{A}_{\tau})$. The smallest ordinal $\sigma$ so that $M\in{\rm Ob}(\mathcal{A}_{\sigma})$ is denoted by $\operatorname{KG-dim}M$. We observe that $\operatorname{KG-dim}0=-1$ and $\operatorname{KG-dim}M\leq 0$ if and only if $\operatorname{ASupp}M\subseteq\textnormal{m-}\operatorname{ASpec}\mathcal{A}$.
We note that any locally noetherian Grothendieck category is semi-noetherian; see [Po, Chap. 5, Theorem 8.5]. To be more precise, If $\mathcal{A}\neq\mathcal{A}_{\tau}$, then $\mathcal{A}/\mathcal{A}_{\tau}$ is also locally noetherian and so it has a nonzero noetherian object $X$. Then $X$ has a maximal subobject $Y$ so that $S=X/Y$ is simple. Therefore, $\sigma=\overline{S}\in\operatorname{ASupp}(\mathcal{A}/\mathcal{A}_{\tau})_{0}$ which is a contradiction by the choice of $\tau$.
Definition 2.13.
Given an ordinal $\sigma\geq 0$, we recall from [MR, GW] that an object $M$ in $\mathcal{A}$ is $\sigma$-critical provided $\operatorname{KG-dim}M=\sigma$ while $\operatorname{KG-dim}M/N<\sigma$ for all nonzero subobjects $N$ of $M$. It is clear that any nonzero subobject of a $\sigma$-critical object is $\sigma$-critical. An object $M$ is said to be critical if it is $\sigma$-critical for some ordinal $\sigma$. It is clear to see that any critical object is monoform.
Lemma 2.14.
Let $M$ be a $\sigma$-critical object in $\mathcal{A}$. Then $\sigma$ is a non-limit ordinal.
Proof.
Assume that $\sigma$ is a limit ordinal. Then there exists some $\rho<\sigma$ such that $t_{\rho}(M)\neq 0$ and so $\operatorname{KG-dim}t_{\rho}(M)\leq\rho$. But $t_{\rho}(M)$ is $\sigma$-critical which is a contradiction. ∎
Lemma 2.15.
Let $\sigma$ be a non-limit ordinal and let $M$ be an object in $\mathcal{A}$. If $F_{\sigma-1}(M)$ is simple, then $M/t_{\sigma-1}(M)$ is $\sigma$-critical.
Proof.
Observe that $F_{\sigma-1}(M)\cong F_{\sigma-1}(M/t_{\sigma-1}(M))$ and so we may assume that $t_{\sigma-1}(M)=0$. Let $N$ be a nonzero subobject of $M$. Then $F_{\sigma-1}(N)$ is nonzero because $t_{\sigma-1}(M)=0$. Since $F_{\sigma-1}(M)$ is simple, we have $F_{\sigma-1}(M/N)=0$ and hence $\operatorname{KG-dim}M/N<\sigma$. On the other hand, by the definition and the fact that $F_{\sigma-1}(M)$ is simple, we have $\operatorname{KG-dim}M=\sigma$. ∎
For every ordinal $\sigma$, the localizing subcategory $\mathcal{A}_{\sigma}$ of $\mathcal{A}$ is generated by critical objects.
Proposition 2.16.
Let $\sigma$ be an ordinal. Then $\mathcal{A}_{\sigma}$ is generated by all $\delta$-critical objects in $\mathcal{A}$ with $\delta\leq\sigma$.
Proof.
If $\sigma$ is a limit ordinal, then $\mathcal{A}_{\sigma}$ is generated by $\bigcup_{\rho<\sigma}\mathcal{A}_{\rho}$ and so we may assume that $\sigma$ is a non-limit ordinal. Let $\mathcal{C}$ be the subclass of all $\delta$-critical objects in $\mathcal{A}$ with $\delta\leq\sigma$ and we have to show that $\mathcal{A}_{\sigma}=\left\langle\mathcal{C}\right\rangle_{\rm loc}$. We prove the claim by transfinite induction on $\sigma$. The case $\sigma=0$ is clear and so we assume that $\sigma>0$. Let $\mathcal{D}$ be the class of $\sigma$-critical objects in $\mathcal{A}$. Then we have the following equalities
$F_{\sigma-1}(\mathcal{A}_{\sigma})=\left\langle F_{\sigma-1}(C)|\hskip 2.84544ptF_{\sigma-1}(C)\hskip 2.84544pt{\rm is\hskip 2.84544ptsimple\hskip 2.84544ptfor\hskip 2.84544pt}C\in\mathcal{A}\right\rangle_{\rm loc}=F_{\sigma-1}(\left\langle\mathcal{A}_{\sigma-1}\cup\mathcal{D}\right\rangle_{\rm loc})=F_{\sigma-1}(\left\langle\mathcal{C}\right\rangle_{\rm loc})$
where the first equality holds by the definition and the second equality holds by [K, Proposition 4.18] and section˜2 and the last equality holds using the induction hypothesis. It now follows from [K, Proposition 4.14] that $\mathcal{A}_{\sigma}=\left\langle\mathcal{C}\right\rangle_{\rm loc}$. ∎
The following lemma is known for modules (for example see [GW, Lemma 15.8]); but since the dimension is slightly different, we reprove it.
Lemma 2.17.
If $M$ is a nonzero object in $\mathcal{A}$ with Krull-Gabriel dimension, then $M$ has a nonzero critical subobject (and so a nonzero monoform subobject)
Proof.
Since ordinals satisfy the descending chain condition, we can choose a nonzero subobject $N$ of $M$ of minimal Krull-Gabriel dimension $\sigma$. Clearly $\sigma$ is non-limit ordinal and $t_{\sigma-1}(N)=0$. Since $F_{\sigma-1}(N)\in(\mathcal{A}/\mathcal{A}_{\sigma-1})_{0}$, it follows from [St, Chap VI, Proposition 2.5] that $F_{\sigma-1}(N)$ contains a simple subobject $S$. Then $N$ contains a subobject $H$ such that $F_{\sigma-1}(H)=S$ by [Po, Chap 4, Corollary 3.10]. Now, section˜2 implies that $H$ is a $\sigma$-critical. ∎
Corollary 2.18.
If $M$ is an object in $\mathcal{A}$ with Krull-Gabriel dimension, then $M$ is loally monoform. In particular, if $\mathcal{A}$ is a semi-noetherian category, then $\mathcal{A}$ is locally monoform.
Proof.
If $M$ has Krull-Gabriel dimension, then every nonzero quotient of $M$ has Krull-Gabriel dimension as well. Thus section˜2 implies that every quotient of $M$ contains a monoform subobject so that $M$ is locally monoform. The second assertion is clear. ∎
Proposition 2.19.
The category $\mathcal{A}$ is semi-noetherian if and only if every nonzero object contains a nonzero critical subobject.
Proof.
If $\mathcal{A}$ is semi-noetherian, by section˜2, every nonzero module has a nonzero critical subobject. Conversely if $\mathcal{A}_{\tau}$ is a proper localizing subcategory of $\mathcal{A}$, then $\mathcal{A}/\mathcal{A}_{\tau}$ contains a nonzero object $Y$. Then there exists a nonzero object $X$ in $\mathcal{A}$ such that $t_{\tau}(X)=0$ and $F_{\tau}(X)=Y$ where $F_{\tau}:\mathcal{A}\rightarrow\mathcal{A}/\mathcal{A}_{\tau}$ is the canonical functor. Hence, by the assumption, $X$ contains a nonzero $\lambda+1$-critical subobject $C$, where $\lambda\geq\tau$. Consequently, by section˜2, we have $\mathcal{A}_{\tau}\neq\mathcal{A}_{\lambda+1}$ which is a contardiction. ∎
3. The atom spectrum of a locally coherent category
A finitely generated object $Y$ in $\mathcal{A}$ is finitely presented if every epimorphism $f:X\rightarrow Y$ in $\mathcal{A}$ with $X$ finitely generated has a finitely generated kernel $\operatorname{Ker}f$. A finitely presented object in $\mathcal{A}$ is coherent if every its finitely generated subobject is finitely presented. We denote by fg-$\mathcal{A}$, fp-$\mathcal{A}$ and coh-$\mathcal{A}$, the full subcategories of $\mathcal{A}$ consisting of finitely generated, finitely presented and coherent objects, respectively.
We recall that a Grothendieck category $\mathcal{A}$ is locally coherent if every object in $\mathcal{A}$ is a direct limit of coherent objects; or equivalently finitely generated subobjects of finitely presented objects are finitely presented. According to [Ro, 2] and [H] a Grothendieck category $\mathcal{A}$ is locally coherent if and only if fp-$\mathcal{A}=$coh-$\mathcal{A}$ is an abelian category. A localizing subcategory $\mathcal{X}$ of $\mathcal{A}$ is said to be of finite type provide that the corresponding right adjoint of the inclusion functor $\mathcal{X}\rightarrow\mathcal{A}$ commutes with direct limits. If $\mathcal{A}$ is a locally noetherian Grothendieck category, then fg-$\mathcal{A}=$noeth-$\mathcal{A}=$fp-$\mathcal{A}=$coh-$\mathcal{A}$ so that $\mathcal{A}$ is locally coherent. In this case, any localizing subcategory $\mathcal{X}$ of $\mathcal{A}$ is of finite type. Throughout this section $\mathcal{A}$ is a locally coherent category.
In our investigation, we use the following lemma due to Krause [Kr, Lemma 1.1].
Lemma 3.1.
An object $X\in\mathcal{A}$ is finitely generated if and only is for any epimorphism $\varphi:Y\rightarrow X$ in $\mathcal{A}$, there is a finitely generated subobject $U$ of $Y$ such that $\varphi(U)=X$
Lemma 3.2.
Let $M$ be an object in $\mathcal{A}$. Then $\operatorname{ASupp}M=\{\alpha\in\operatorname{ASpec}\mathcal{A}|\hskip 2.84544pt\operatorname{Hom}(M,E(\alpha))\neq 0\}$.
Proof.
Given $\alpha\in\operatorname{ASupp}M$, there exist subobjects $K\subset L\subseteq M$ such that $H=L/K$ is a monoform object with $\alpha=\overline{H}$. Since $\operatorname{Hom}(H,E(\alpha))\neq 0$, we have $\operatorname{Hom}(L,E(\alpha))\neq 0$ and consequently $\operatorname{Hom}(M,E(\alpha))\neq 0$. ∎
For every subcategory $\mathcal{S}$ of $\mathcal{A}$, we denote by $\overset{\rightarrow}{\mathcal{S}}$, the full subcategory of $\mathcal{A}$ consisting of direct limits $\underset{\rightarrow}{\rm lim}X_{i}$ with $X_{i}\in\mathcal{S}$ for each $i$. For $\mathcal{S}\subset\operatorname{fp-}\mathcal{A}$, we denote by $\surd{\mathcal{S}}$, the smallest Serre subcategory of $\operatorname{fp-}\mathcal{A}$ containing $\mathcal{S}$. The following lemma establishes another topology on $\operatorname{ASpec}\mathcal{A}$.
Lemma 3.3.
The set $\{\operatorname{ASupp}M|\hskip 2.84544ptM\in\operatorname{fp-}\mathcal{A}\}$ forms a basis of open subsets for a topology on $\operatorname{ASpec}\mathcal{A}$. Furthermore, if $M\in\operatorname{fp-}\mathcal{A}$, then $\operatorname{ASupp}M=\emptyset$ if and only $M=0$.
Proof.
Since $\mathcal{A}$ is locally coherent, it is clear that for every $\alpha\in\operatorname{ASpec}\mathcal{A}$, there exists a finitely presented object $M$ in $\mathcal{A}$ such that $\alpha\in\operatorname{ASupp}M$. If $M_{1}$ and $M_{2}$ are finitely presented objects in $\mathcal{A}$ and $\alpha\in\operatorname{ASupp}M_{1}\cap\operatorname{ASupp}M_{2}$, then by section˜3, there exists a nonzero morphism $f_{i}:M_{i}\rightarrow E(\alpha)$ for $i=1,2$. Since $E(\alpha)$ is uniform, $\operatorname{Im}f_{1}\cap\operatorname{Im}f_{2}$ is a nonzero subobject of $E(\alpha)$ and so it contains a nonzero finitely generated subobject $X$ as $\mathcal{A}$ is locally coherent. Using the pull-back diagram and section˜3, there exists a finitely presented subobject $L_{i}$ of $M_{i}$ such that $f_{i}(L_{i})=X$ for $i=1,2$. This implies that $X\in\overset{\longrightarrow}{\surd{M_{1}}}\cap\overset{\longrightarrow}{\surd{M_{2}}}$. By virtue of [H, Proposition 2.3], the morphism $f_{1}:L_{1}\rightarrow X$ factors through a quotient $N$ of $L_{1}$ which lies in $\surd{M_{2}}$. Therefore $N$ is finitely presented and $\alpha\in\operatorname{ASupp}N\subset\operatorname{ASupp}M_{1}\cap\operatorname{ASupp}M_{2}$. To prove the second claim, every nonzero finitely presented object $M$ in $\mathcal{A}$ has a simple quotient object $S$ so that $\alpha=\overline{S}\in\operatorname{ASupp}M$. ∎
To prevent any misunderstanding, we use the symbol $\operatorname{ZASpec}\mathcal{A}$ instead of $\operatorname{ASpec}\mathcal{A}$ with the new topology based on section˜3. We notice that m-$\operatorname{ASpec}\mathcal{A}$ is a dense subset of $\operatorname{ZASpec}\mathcal{A}$. Because if $M$ is a finitely presented object in $\mathcal{A}$, it contains a maximal subobject $N$ so that the maximal atom $\overline{M/N}\in\operatorname{ASupp}M$. We also remark that if $M$ is a finitely presented object in $\mathcal{A}$, then $\operatorname{ASupp}K$ is an open subset of $\operatorname{ZASpec}\mathcal{A}$ for any subobject $K$ of $M$. Because $K=\bigcup_{i}K_{i}$ is the direct union of its finitely generated subobjects. Since $M$ is finitely presented, each $K_{i}$ is finitely presented so that $\operatorname{ASupp}K=\bigcup_{i}\operatorname{ASupp}K_{i}$ is an open subset of $\operatorname{ZASpec}\mathcal{A}$.
Theorem 3.4.
The map $\mathcal{X}\mapsto\operatorname{ASupp}\mathcal{X}$ provides a bijective correspondence between the lattice of Serre subcategories of $\operatorname{fp-}\mathcal{A}$ and the lattice of open subsets of $\operatorname{ZASpec}\mathcal{A}$. The inverse map is given by $\mathcal{U}\mapsto\operatorname{ASupp}^{-1}\mathcal{U}\cap\operatorname{fp-}\mathcal{A}$.
Proof.
Assume that $\mathcal{X}$ is a Serre subcategory of $\operatorname{fp-}\mathcal{A}$ and $\mathcal{U}$ is an open subset of $\operatorname{ZASpec}\mathcal{A}$. It is clear that $\operatorname{ASupp}\mathcal{X}$ is an open subset of $\operatorname{ZASpec}\mathcal{A}$ and $\operatorname{ASupp}^{-1}\mathcal{U}\cap\operatorname{fp-}\mathcal{A}$ is a Serre subcategory of $\operatorname{fp-}\mathcal{A}$. In order to prove $\operatorname{ASupp}^{-1}(\operatorname{ASupp}\mathcal{X})\cap\operatorname{fp-}\mathcal{A}=\mathcal{X}$, it suffices to show that $\operatorname{ASupp}^{-1}(\operatorname{ASupp}\mathcal{X})\cap\operatorname{fp-}\mathcal{A}\subset\mathcal{X}$. Given $M\in\operatorname{ASupp}^{-1}(\operatorname{ASupp}\mathcal{X})\cap\operatorname{fp-}\mathcal{A}$, we have $\operatorname{ASupp}M\subset\operatorname{ASupp}\mathcal{X}$ and so $\operatorname{ASupp}M\subset\operatorname{ASupp}\overset{\longrightarrow}{\mathcal{X}}$. It follows from [H, Proposition 2.15] that $F(M)$ is a finitely presented object in $\mathcal{A}/\overset{\longrightarrow}{\mathcal{X}}$, where $F:\mathcal{A}\rightarrow\mathcal{A}/\overset{\longrightarrow}{\mathcal{X}}$ is the canonical functor. Moreover, $\operatorname{ASupp}F(M)=\operatorname{ASupp}M\setminus\operatorname{ASupp}\mathcal{X}=\emptyset$ by [K, Proposition 5.6]; and hence $F(M)=0$ by section˜3. Therefore, $M\in\overset{\longrightarrow}{\mathcal{X}}$ and so $M=\underset{\rightarrow}{\rm lim}M_{i}$ is the direct limit of objects $M_{i}$ in $\mathcal{X}$. Since $M$ is finitely presented, it is a direct summand of some $M_{i}$ so that $M\in\mathcal{X}$. The fact that $\operatorname{ASupp}(\operatorname{ASupp}^{-1}\mathcal{U})=\mathcal{U}$ is straightforward. ∎
For a localizing subcategory $\mathcal{X}$ of finite type of $\mathcal{A}$, we define $\operatorname{ASupp}\mathcal{X}=\operatorname{ASupp}(\mathcal{X}\cap\operatorname{fp-}\mathcal{A})$ which is an open subset of $\operatorname{ZASpec}\mathcal{A}$.
Corollary 3.5.
The map $\mathcal{U}\mapsto\overrightarrow{\operatorname{ASupp}^{-1}\mathcal{U}\cap\operatorname{fp-}\mathcal{A}}$ provides a bijective correspondence between the lattice of open subsets of $\operatorname{ZASpec}\mathcal{A}$ and the lattice of localizing subcategories of finite type of $\mathcal{A}$. The inverse map is given by $\mathcal{X}\mapsto\operatorname{ASupp}\mathcal{X}$.
Proof.
Given an open subset $\mathcal{U}$ of $\operatorname{ZASpec}\mathcal{A}$, by [Kr, Theorem 2.8], $\overrightarrow{\operatorname{ASupp}^{-1}\mathcal{U}\cap\operatorname{fp-}\mathcal{A}}$ is a localizing subcategory of finite type of $\mathcal{A}$. It is easy to see that $\operatorname{ASupp}(\overrightarrow{\operatorname{ASupp}^{-1}\mathcal{U}\cap\operatorname{fp-}\mathcal{A}})=\mathcal{U}$. Also, for every localizing subcategory $\mathcal{X}$ of finite type of $\mathcal{A}$, it follows from section˜3 and [Kr, Lemma 2.3] that
$\overrightarrow{\operatorname{ASupp}^{-1}(\operatorname{ASupp}\mathcal{X})\cap\operatorname{fp-}\mathcal{A}}=\overrightarrow{\operatorname{ASupp}^{-1}(\operatorname{ASupp}(\mathcal{X}\cap\operatorname{fp-}\mathcal{A}))\cap\operatorname{fp-}\mathcal{A}}=\overrightarrow{\mathcal{X}\cap\operatorname{fp-}\mathcal{A}}=\mathcal{X}.$
The following corollary provides a characterization for localizing subcategories of finite type of $\mathcal{A}$ in terms of atoms.
Corollary 3.6.
Let $\mathcal{X}$ be a localizing subcategory of $\mathcal{A}$. Then $\mathcal{X}$ is of finite type if and only if $\operatorname{ASupp}\mathcal{X}$ is an open subset of $\operatorname{ZASpec}\mathcal{A}$.
Proof.
If $\operatorname{ASupp}\mathcal{X}$ is an open subset of $\operatorname{ZASpec}\mathcal{A}$, then section˜3 implies that $\mathcal{X}=\overrightarrow{\operatorname{ASupp}^{-1}(\operatorname{ASupp}\mathcal{X})\cap\operatorname{fp-}\mathcal{A}}$ is of finite type of $\mathcal{A}$. The converse is clear. ∎
An open cover of a topological space $\mathcal{U}$ is a family $\{\mathcal{U}_{\lambda}|\hskip 2.84544pt\lambda\in\varLambda\}$ of open subsets of $\mathcal{U}$ satisfying $\mathcal{U}=\bigcup_{\lambda\in\varLambda}U_{\lambda}$. A topological space $\mathcal{U}$ is quasi-compact if for any open cover $\{\mathcal{U}_{i}\}_{\lambda\in\varLambda}$ of $\mathcal{U}$, there exists a finite subset $\varGamma$ of $\varLambda$ such that $\mathcal{V}=\bigcup_{\gamma\in\varGamma}\mathcal{U}_{\gamma}$.
Proposition 3.7.
An open subset $\varPhi$ of $\operatorname{ZASpec}\mathcal{A}$ is a quasi-compact if and only if there exists an object $M$ in $\operatorname{fp-}\mathcal{A}$ such that $\varPhi=\operatorname{ASupp}M$.
Proof.
Let $\varPhi$ be quasi-compact. Since $\varPhi$ is open, by definition, we have
$\varPhi=\bigcup_{M\in\operatorname{ASupp}^{-1}\varPhi\cap\operatorname{fp-}\mathcal{A}}\operatorname{ASupp}M.$
Then there exists $M_{1},\dots,M_{n}\in\operatorname{ASupp}^{-1}\varPhi$ such that $\varPhi=\bigcup_{i=1}^{n}\operatorname{ASupp}M_{i}=\operatorname{ASupp}(\bigoplus_{i=1}^{n}M_{i})$. Then the result follows as $\bigoplus_{i=1}^{n}M_{i}$ is finitely presented. Conversely, assume that $M$ is a finitely presented object in $\mathcal{A}$ and $\{\varPhi_{\lambda}|\hskip 2.84544pt\lambda\in\varLambda\}$ is an open cover of $\operatorname{ASupp}M$. It is clear that
$\bigcup_{\lambda\in\varLambda}\varPhi_{\lambda}=\operatorname{ASupp}\sqrt{\langle\operatorname{ASupp}^{-1}\varPhi_{\lambda}\cap\operatorname{fp-}\mathcal{A}|\hskip 2.84544pt\lambda\in\varLambda\rangle}$
so that
$\operatorname{ASupp}M=\bigcup_{\lambda\in\varLambda}\varPhi_{\lambda}=\operatorname{ASupp}\sqrt{\langle\operatorname{ASupp}^{-1}\varPhi_{\lambda}\cap\operatorname{fp-}\mathcal{A}|\hskip 2.84544pt\lambda\in\varLambda\rangle}$
$\subseteq\operatorname{ASupp}\overrightarrow{\sqrt{\langle\operatorname{ASupp}^{-1}\varPhi_{\lambda}\cap\operatorname{fp-}\mathcal{A}|\hskip 2.84544pt\lambda\in\varLambda\rangle}}.$
Hence section˜3 implies that $M\in\overrightarrow{\sqrt{\langle\operatorname{ASupp}^{-1}\varPhi_{\lambda}\cap\operatorname{fp-}\mathcal{A}|\hskip 2.84544pt\lambda\in\varLambda\rangle}}$. Since $M$ is finitely presented, it is the direct summand of some object in $\sqrt{\langle\operatorname{ASupp}^{-1}\varPhi_{\lambda}\cap\operatorname{fp-}\mathcal{A}|\hskip 2.84544pt\lambda\in\varLambda\rangle}$ and so $M\in\sqrt{\langle\operatorname{ASupp}^{-1}\varPhi_{\lambda}\cap\operatorname{fp-}\mathcal{A}|\hskip 2.84544pt\lambda\in\varLambda\rangle}$. Therefore $M$ has a finite filtration of finitely presented subobjects $0=M_{0}\subset M_{1}\subset M_{2}\subset\dots\subset M_{n}=M$ such that $M_{i}/M_{i-1}$ is a subquotient of some finitely presented object $N_{\lambda_{i}}\in\operatorname{ASupp}^{-1}\varPhi_{\lambda_{i}}$ for $1\leq i\leq n$. Therefore $\operatorname{ASupp}M=\bigcup_{i=1}^{n}\varPhi_{\lambda_{i}}$
For a locally coherent Grothendieck category $\mathcal{A}$, the subcategory $\operatorname{fp-}\mathcal{A}$ is abelian. Thus, the atom spectrum of $\operatorname{fp-}\mathcal{A}$ can be investigated independently. To prevent any misunderstanding, any atom in $\operatorname{ASpec}\operatorname{fp-}\mathcal{A}$ is denoted by $\alpha_{f}$ and for every object $M$ in $\operatorname{fp-}\mathcal{A}$, we use the symbol $\operatorname{fASupp}M$ for atom support of $M$ in $\operatorname{ASpec}\operatorname{fp-}\mathcal{A}$ instead of $\operatorname{ASupp}M$. Similarly we use the symbol $\operatorname{fAAss}M$ instead of $\operatorname{AAss}M$.
Lemma 3.8.
Any monoform object in $\operatorname{fp-}\mathcal{A}$ is a monoform object in $\mathcal{A}$.
Proof.
Let $H$ be a monoform object in $\operatorname{fp-}\mathcal{A}$. If $H$ is not monoform object in $\mathcal{A}$, then there exists a subobject $N$ of $H$ and a nonzero finitely presented subobject $X$ of $H/N$ isomorphic to a subobject of $H$. Since $\operatorname{fp-}\mathcal{A}$ is abelian, using a pullback diagram and section˜3, there exist finitely generated (and so finitely presented) subobjects $C\subset D$ of $H$ such that $X\cong D/C$. This implies that $H$ and $H/C$ has a common nonzero subobject which contradicts the fact that $H$ is a monoform object of $\operatorname{fp-}\mathcal{A}$. ∎
Corollary 3.9.
Let $M$ be a finitely presented object in $\mathcal{A}$. Then $\operatorname{fAAss}M=(\operatorname{AAss}M)_{f}$.
Proof.
If $\alpha\in\operatorname{AAss}M$, then $M$ contains a monoform subobject $H$ in $\mathcal{A}$ such that $\alpha=\overline{H}$. Since $\mathcal{A}$ is locally coherent, we may assume that $H$ is finitely generated. Hence $H$ is finitely presented because $M$ is finitely presented. Moreover, it is clear that $H$ is a monoform object of $\operatorname{fp-}\mathcal{A}$ and consequently $\alpha_{f}\in\operatorname{fAAss}M$. To prove the other side, if $\alpha_{f}\in\operatorname{fAAss}M$, then $M$ contains a monoform subobject $H$ of $\operatorname{fp-}\mathcal{A}$ such that $\alpha=\overline{H}$. It follows from section˜3 that $H$ is a monoform object in $\mathcal{A}$ so that $\alpha\in\operatorname{AAss}M$. Conversely if $\alpha\in\operatorname{AAss}M$, since $M$ is finitely presented, it contains a monoform subobject $H$ such that $\overline{H}=\alpha$. ∎
The atom spectrum $\operatorname{ASpec}\operatorname{fp-}\mathcal{A}$ can be considered as a subspace of $\operatorname{ZASpec}\mathcal{A}$.
Proposition 3.10.
The map $\alpha_{f}\stackrel{{\scriptstyle g}}{{\mapsto}}\alpha$ embeds $\operatorname{ASpec}\operatorname{fp-}\mathcal{A}$ as a subspace of $\operatorname{ZASpec}\mathcal{A}$. Moreover, if $\operatorname{fp-}\mathcal{A}$ is locally monoform, then $\operatorname{ASpec}\operatorname{fp-}\mathcal{A}$ is dense in $\operatorname{ZASpec}\mathcal{A}$.
Proof.
Given $\alpha_{f}\in\operatorname{ASpec}\operatorname{fp-}\mathcal{A}$, there exists a monoform object $H$ of the abelian category $\operatorname{fp-}\mathcal{A}$ such that $\alpha_{f}=\overline{H}$. Thus $\alpha=\overline{H}\in\operatorname{ZASpec}\mathcal{A}$ and so identifying $\alpha_{f}$ with $\alpha$, we show that $\operatorname{ASpec}\operatorname{fp-}\mathcal{A}$ is a topological subspace of $\operatorname{ZASpec}\mathcal{A}$. It is clear that $g$ is injective and we show that $g$ is continuous. For any finitely presented object $M$ in $\mathcal{A}$, we prove that $g^{-1}(\operatorname{ASupp}M)=\operatorname{fASupp}M$. The above argument indicates that $\operatorname{fASupp}M\subset g^{-1}(\operatorname{ASupp}M)$. To prove the converse, assume that $\alpha_{f}\in g^{-1}(\operatorname{ASupp}M)$. By section˜3, there exists a finitely presented monoform object $H$ in $\mathcal{A}$ such that $\alpha=\overline{H}$ and $H$ is a subquotient of $M$. Using section˜3, we can choose such $H$ such that $H=L/K$, where $L$ is a finitely presented subobject of $M$. Since $\operatorname{fp-}\mathcal{A}$ is abelian, we deduce that $K$ is finitely presented; and hence $\alpha_{f}\in\operatorname{fASupp}M$. To prove the second claim, the assumption and section˜3 imply that every finitely presented object $M$ contains a finitely presented monoform subobject $H$ in $\mathcal{A}$. Then $\overline{H}=\alpha_{f}\in\operatorname{ASupp}M\cap\operatorname{ASpec}\operatorname{fp-}\mathcal{A}$. ∎
It is known as Cohen’s theorem that if prime ideals of a commutative ring $A$ are finitely generated, then $A$ is noetherian. In the following theorem we extend this theorem for locally coherent categories.
Theorem 3.11.
If $\operatorname{ASpec}\operatorname{fp-}\mathcal{A}=\operatorname{ASpec}\mathcal{A}$ (i.e. if g in section˜3 is surjective), then $\mathcal{A}$ is locally noetherian.
Proof.
Given a finitely presented object $M$ in $\mathcal{A}$, it suffices to show that $M$ is noetherian. Otherwise, $M$ contains a subobject which is not finitely generated. It follows from [St, Chap IV, Proposition 6.6] that the class of subobjects of $M$ is a set. Then the set of subobjects of $M$
$\{N\subset M|\hskip 2.84544ptN\hskip 2.84544ptis\hskip 2.84544pt{\rm not}\hskip 2.84544pt{\rm finitely}\hskip 2.84544pt{\rm generated}\}$
has a maximal element $K$ by Zorn’s lemma. We claim that $M/K$ is a monoform object in $\mathcal{A}$. Otherwise, there exists a subobject $L$ of $M$ containing $K$ such that $M/K$ and $M/L$ has a common nonzero subobject $X$. The maximality $K$ implies that $L$ is finitely generated so that $M/L$ is finitely presented. Then we may assume that $X$ is finitely presented. Hence there exists a subobject $K_{1}$ of $M$ containig $K$ such that $X\cong K_{1}/K$ is finitely presented. Again maximality implies that $K_{1}$ is finitely generated and so is finitely presented. Since $\operatorname{fp-}\mathcal{A}$ is abelian, we deduce that $K$ is finitely presented which is a contradiction. Therefore $\overline{M/K}\in\operatorname{ASpec}\mathcal{A}$ and so the assumption implies that $M/K$ contains a finitely presented monoform subobject $Y/K$. Again by maximality of $K$, the subobject $Y$ is finitely presented so that $K$ is finitely presented which is a contradiction. ∎
In view of section˜3, for any object $M$ in $\mathcal{A}$, we define the following subsets of $\operatorname{ASpec}\operatorname{fp-}\mathcal{A}$:
$\operatorname{fASupp}M=\{\alpha_{f}=\overline{H}\in\operatorname{ASpec}\operatorname{fp-}\mathcal{A}|\hskip 2.84544pt\alpha=\overline{H}\in\operatorname{ASupp}M\};$
$\operatorname{fAAss}M=\{\alpha_{f}=\overline{H}\in\operatorname{ASpec}\operatorname{fp-}\mathcal{A}|\hskip 2.84544pt\alpha=\overline{H}\in\operatorname{AAss}M\}.$
For any subcategory $\mathcal{X}$ of $\mathcal{A}$ , we define $\operatorname{fASupp}\mathcal{X}=\bigcup_{M\in\mathcal{X}}\operatorname{fASupp}M$. In the following theorem, we provide a sufficient condition under which a localizing subcategory of $\mathcal{A}$ is of finite type.
Theorem 3.12.
Any localizing subcategory $\mathcal{X}$ of $\mathcal{A}$ with $\operatorname{fASupp}\mathcal{X}=\operatorname{ASupp}\mathcal{X}$ is of finite type.
Proof.
By [Kr, Theorem 2.8], it suffices to show that $\mathcal{X}=\overrightarrow{\mathcal{X}\cap\operatorname{fp-}\mathcal{A}}$. The inclusion $\overrightarrow{\mathcal{X}\cap\operatorname{fp-}\mathcal{A}}\subset\mathcal{X}$ is clear. For the other side, since $\mathcal{A}$ is a locally coherent, $M=\bigcup_{i}M_{i}$ is the direct union of its finitely generated subobjects. Then it suffices to show that each $M_{i}$ is finitely presented. Hence, we may assume that $M$ is a finitely generated object in $\mathcal{X}$ and we have to prove that $M$ is finitely presented. Since $\mathcal{A}$ is locally coherent, there exists a finitely presented object $N$ in $\mathcal{A}$ and a subobject $L$ of $N$ such that $M=N/L$. If $L$ is not finitely generated, the set
$\{X\subset N|\hskip 2.84544ptL\subset X\hskip 2.84544pt{\rm and}\hskip 2.84544ptX\hskip 2.84544ptis\hskip 2.84544pt{\rm not}\hskip 2.84544pt{\rm finitely}\hskip 2.84544pt{\rm generated}\}$
has a maximal element $K$ by Zorn’s lemma. By a similar proof as given in section˜3, we can deduce that $N/K$ is a monoform object in $\mathcal{A}$. Thus $\overline{N/K}\in\operatorname{ASupp}M$ and the assumption implies that $N/K$ contains a finitely presented subobject $Y$. Therefore by the maximality of $K$, NN has a finitely presented subobject $Z$ containing $K$ such that $Y=Z/K$. Since $\operatorname{fp-}\mathcal{A}$ is abelian, we conclude that $K$ is finitely presented which is a contradiction. Hence $L$ is finitely generated so that $M$ is finitely presented. ∎
4. Ziegler spectrum of a locally coherent category
Throughout this section $\mathcal{A}$ is a locally coherent. The collection of the isomorphism classes of indecomposable injective objects in $\mathcal{A}$, denoted by $\operatorname{Zg}\mathcal{A}$, forms a set. This is because every indecomposable injective object is the injective envelope of some quotient of an element of the generating set of $\mathcal{A}$.
Definition 4.1.
The Ziegler support $\operatorname{\mathcal{O}}(M)$ of an object $M$ in $\mathcal{A}$ is a subset of $\operatorname{Zg}\mathcal{A}$, that is
$\operatorname{\mathcal{O}}(M)=\{I\in\operatorname{Zg}\mathcal{A}|\hskip 2.84544pt\operatorname{Hom}(M,I)\neq 0\}.$
For any exact sequence $0\rightarrow N\rightarrow M\rightarrow K\rightarrow 0$ of objects in $\mathcal{A}$, we have
$\operatorname{\mathcal{O}}(M)=\operatorname{\mathcal{O}}(N)\cup\operatorname{\mathcal{O}}(K).$
For any subcategory $\mathcal{X}$ of $\mathcal{A}$, we define $\operatorname{\mathcal{O}}(\mathcal{X})=\bigcup_{M\in\mathcal{X}}\operatorname{\mathcal{O}}(M)$ and for any subset $\mathcal{U}$ of $\operatorname{Zg}\mathcal{A}$, we define $\operatorname{\mathcal{O}}^{-1}(\mathcal{U})=\{M\in\mathcal{A}|\hskip 2.84544pt\operatorname{\mathcal{O}}(M)\subset\mathcal{U}\}$.
Herzog [H] proved that the collection $\{\mathcal{O}(M)|\hskip 2.84544ptM\in\operatorname{fp-}\mathcal{A}\}$ satisfies the axioms for a basis of open subsets defining a topology on $\operatorname{Zg}\mathcal{A}$. The resulting topological space is called the Ziegler spectrum of $\mathcal{A}$.
For $\alpha\in\operatorname{ASpec}\mathcal{A}$ and monoform objects $H_{1}$ and $H_{2}$ in $\mathcal{A}$ satisfying $\alpha=\overline{H_{1}}=\overline{H_{2}}$, we have $E(H_{1})=E(H_{2})$. The isomorphism class of all such $E(H)$ such that $\alpha=\overline{H}$ is denoted by $E(\alpha)$. We observe that $E(\alpha)$ is an indecomposable injective object in $\mathcal{A}$. Because if $E(\alpha)=E(H)$ for some monoform object $H$ in $\mathcal{A}$ and $E(\alpha)=E_{1}\oplus E_{2}$, then $E_{1}\cap H$ and $E_{2}\cap H$ are nonzero monoform subobjecs of $H$. Thus $E_{1}\cap H\cap E_{2}\cap H$ is nonzero which is a contradiction. Then we may identify $\operatorname{ZASpec}\mathcal{A}$ with a subset of $\operatorname{Zg}\mathcal{A}$ via the map $\alpha\mapsto E(\alpha)$. Furthermore, we have the following lemma.
Lemma 4.2.
There exists a continuous injective map $\theta:\operatorname{ZASpec}\mathcal{A}\rightarrow\operatorname{Zg}\mathcal{A}$, given by $\alpha\mapsto E(\alpha)$. In particular, $\operatorname{ZASpec}\mathcal{A}$ is a dense subset of $\operatorname{Zg}\mathcal{A}$.
Proof.
The injectivity of $\theta$ is clear. For every $M\in\operatorname{fp-}\mathcal{A}$, it follows from section˜3 that $\theta^{-1}(\mathcal{O}(M))=\operatorname{ASupp}M$. Then $\theta^{-1}(\mathcal{O})$ is an open subset of $\operatorname{ZASpec}\mathcal{A}$ for any open subset $\mathcal{O}$ of $\operatorname{Zg}\mathcal{A}$. To prove the second claim, every finitely presented object $M$ has a maximal subobject $N$ so that $M/N$ is a monoform object in $\mathcal{A}$. If we consider $\alpha=\overline{M/N}$, it is clear that $\operatorname{Hom}(M,E(\alpha))\neq 0$. ∎
A category $\mathcal{A}$ is said to be locally uniform (or locally coirreducible by [Po, p. 330]) if every nonzero object in $\mathcal{A}$ has a uniform subobject. The following proposition establishes a characterization of locally monoform objects.
Proposition 4.3.
A category $\mathcal{A}$ is locally monoform if and only if $\mathcal{A}$ is locally uniform and the map $\theta:\operatorname{ZASpec}\mathcal{A}\rightarrow\operatorname{Zg}\mathcal{A}$, given by $\alpha\mapsto E(\alpha)$ is a homeomorphism.
Proof.
Let $\mathcal{A}$ be locally monoform. The it is clear that $\mathcal{A}$ is locally uniform. Let $E$ be an indecomposable injective object in $\mathcal{A}$. By the assumption, the object $E$ contains a monoform subobject $H$ and so $E=E(\alpha)$, where $\alpha=\overline{H}$. This implies that $f$ is surjective. Therefore, it follows from section˜4 that $f$ is a homeomorphism. Conversely, let $M$ be an object in $\mathcal{A}$. Since $\mathcal{A}$ is locally uniform, $M$ contains a uniform subobject $U$ so that $E(U)$ is an injective indecomposable. Since $f$ is surjective, there exists a monoform object $\alpha\in\operatorname{ASpec}\mathcal{A}$ such that $E(\alpha)=E(U)$. Then there exists a monoform object $H$ in $\mathcal{A}$ such that $\alpha=\overline{H}$ and $E(H)=E(U)$ so that $H\cap U$ is a monoform subobject of $M$. ∎
For a commutative ring $A$, We have the following easy lemma.
Lemma 4.4.
Let $\mathfrak{p}$ be an ideal of $A$. Then $A/\mathfrak{p}$ is a monoform $A$-module if and only if $\mathfrak{p}$ is a prime ideal of $A$.
Proof.
Clear. ∎
We recall that a topological space $X$ is said to be $T_{0}$-space (or Kolmogorov) if for any distinct points $x,y$ of $X$, there exists an open subset of $X$ containing exactly one of them. The following example [GP], pointed out by T. Kucera, shows that a locally coherent category $\mathcal{A}$ need not be locally monoform and also $\operatorname{Zg}\mathcal{A}$ need not be $T_{0}$-space.
Example 4.5.
Let $A=k[X_{n}(n\in\omega)]$ be a polynomial ring over a field $k$ in infinitely many commuting indeterminates. Then $A$ is clearly a commutative coherent ring. Let $I=(X_{n}^{n+1}:n\in\omega)$. Then $I$ is not prime as $X_{1}\notin I$ but $X_{1}^{2}\in I$. According to [P], the injective $A$-module $E(A/I)$ is indecomposable and it does not have the form $E(A/\mathfrak{p})$ for any prime ideal $\mathfrak{p}$ of $A$ by [P, 9.1]. In section˜4, if $f$ is surjective, then there exists $\alpha\in\operatorname{ASpec}\operatorname{Mod}A$ such that $E(A/I)=E(\alpha)$. Hence there exists a cyclic monoform $A$-module $H$ such that $E(H)=E(A/I)$. But section˜4 implies that $H=A/\mathfrak{p}$ for some prime ideal $\mathfrak{p}$ of $A$ which is a contradiction. Therefore $\operatorname{Mod}A$ is not locally monoform by section˜4. Furthermore, since $E(A/I)$ does not have the form $E(A/\mathfrak{p})$ for any prime ideal, by [GP, Theorem 1.4], there exists a prime ideal $\mathfrak{p}$ of $A$ such that $\varLambda(E(A/I))=\varLambda(E(A/\mathfrak{p}))$, where for any injective indecomposable $E$, the set $\varLambda(E)$ denotes the intersection of all open subsets of $\operatorname{Zg}\operatorname{Mod}A$ containing $E$. But this implies that $\operatorname{Zg}\operatorname{Mod}A$ is not $T_{0}$-space.
The following proposition provide a sufficient condition under which a locally coherent category is semi-noetherian (and so locally monoform).
Proposition 4.6.
If $\mathcal{A}_{\tau}$ is a localizing subcategory of finite type of $\mathcal{A}$, then $\mathcal{A}$ is semi-noetherian.
Proof.
If $\mathcal{A}$ is not semi noetherian, then $\mathcal{A}_{\tau}$ is a proper subcategory of $\mathcal{A}$. It follows from [H, Theorem 2.16] that $\mathcal{A}/\mathcal{A}_{\tau}$ is locally coherent so that it contains a simple object $S$. Hence there exists an object $X\in\mathcal{A}_{\tau+1}\setminus\mathcal{A}_{\tau}$ such that $F_{\tau}(X)=S$ which is a contradiction. ∎
Krause [Kr] has constructed a topology on $\operatorname{Zg}\mathcal{A}$ in which for a subset $\mathcal{U}$ of $\operatorname{Zg}\mathcal{A}$, the closure of $\mathcal{U}$ is defined as $\overline{\mathcal{U}}=\left\langle{}^{\bot}\mathcal{U}\cap\operatorname{fp-}\mathcal{A}\right\rangle^{\bot}$. The subsets $\mathcal{U}$ of $\operatorname{Zg}\mathcal{A}$ satisfying $\mathcal{U}=\overline{\mathcal{U}}$ form the closed subsets of $\operatorname{Zg}\mathcal{A}.$ The following proposition shows that the Ziegler topology and the topology defined by Krause are identical.
Lemma 4.7.
The Ziegler topology and Krause topology on $\operatorname{Zg}(\mathcal{A})$ are the same.
Proof.
We show that $\operatorname{Zg}\mathcal{A}$ and $\operatorname{Sp}\mathcal{A}$ have the same open subsets. Given an open subset $\mathcal{O}$ of $\operatorname{Zg}(\mathcal{A})$, it suffices to show that $\left\langle{}^{\bot}\mathcal{O}^{c}\cap\operatorname{fp-}\mathcal{A}\right\rangle^{\bot}=\mathcal{O}^{c}$ and so $\mathcal{O}^{c}$ will be a closed subset of $\operatorname{Sp}\mathcal{A}$, where $\mathcal{O}^{c}=\operatorname{Sp}\mathcal{A}\setminus\mathcal{O}$. If $I\in\mathcal{O}^{c}$, then it is clear that $\operatorname{Hom}(^{\bot}\mathcal{O}^{c}\cap\operatorname{fp-}\mathcal{A},I)=0$ and so $I\in\left\langle{}^{\bot}\mathcal{O}^{c}\cap\operatorname{fp-}\mathcal{A}\right\rangle^{\bot}$. Conversely, if $I\in\left\langle{}^{\bot}\mathcal{O}^{c}\cap\operatorname{fp-}\mathcal{A}\right\rangle^{\bot}\setminus\mathcal{O}^{c}$, there exists $M\in\operatorname{fp-}\mathcal{A}$ such that $I\in\mathcal{O}(M)\subseteq\mathcal{O}$; and hence $M\notin^{\bot}\mathcal{O}^{c}\cap\operatorname{fp-}\mathcal{A}$. Then $\operatorname{Hom}(M,\mathcal{O}^{c})\neq 0$ so that there exists $J\in\mathcal{O}^{c}$ such that $\operatorname{Hom}(M,J)\neq 0$. But this implies that $J\in\mathcal{O}(M)\subseteq\mathcal{O}$ which is a contradiction. Now suppose that $\mathcal{O}$ is an open subset of $\operatorname{Sp}\mathcal{A}$ and so $\mathcal{O}^{c}=\left\langle{}^{\bot}\mathcal{O}^{c}\cap\operatorname{fp-}\mathcal{A}\right\rangle^{\bot}$. We now show that $\mathcal{O}$ is an open subset of $\operatorname{Zg}\mathcal{A}$. Given $I\in\mathcal{O}$, we have $\operatorname{Hom}(^{\bot}\mathcal{O}^{c}\cap\operatorname{fp-}\mathcal{A},I)\neq 0$ and so there exists $M\in^{\bot}\mathcal{O}^{c}\cap\operatorname{fp-}\mathcal{A}$ such that $\operatorname{Hom}(M,I)\neq 0$. Thus $I\in\mathcal{O}(M)$ and $\operatorname{Hom}(M,\mathcal{O}^{c})=0$. For every $J\in\mathcal{O}(M)$, we have $\operatorname{Hom}(M,J)\neq 0$ which implies that $J\in\mathcal{O}$. Therefore, $\mathcal{O}(M)\subset\mathcal{O}$ ; and consequently $\mathcal{O}$ is an open subset of $\operatorname{Zg}\mathcal{A}$. ∎
The localizing subcategories of finite type of $\mathcal{A}$ can be classifies via their Ziegler supports which are open subsets of $\operatorname{Zg}\mathcal{A}$.
Proposition 4.8.
The map $\mathcal{X}\mapsto\operatorname{\mathcal{O}}(\mathcal{X})$ provides a bijective correspondence between the lattice of localizing subcategories of finite type of $\mathcal{A}$ and the lattice of open subsets of $\operatorname{Zg}\mathcal{A}$. The inverse map is $\mathcal{U}\mapsto\operatorname{\mathcal{O}}^{-1}(\mathcal{U})$.
Proof.
Given a localizing subcategory $\mathcal{X}$ of finite type of $\mathcal{A}$, by [Kr, Lemma 2.3], we have $\mathcal{X}=\overset{\rightarrow}{\mathcal{S}}$, where $\mathcal{S}=\mathcal{X}\cap\operatorname{fp-}\mathcal{A}$. Then $\operatorname{\mathcal{O}}(\mathcal{X})=\bigcup_{M\in\mathcal{S}}\operatorname{\mathcal{O}}(M)$ that is an open subset of $\operatorname{Zg}\mathcal{A}$. Given an open subset $\mathcal{U}$ of $\operatorname{Zg}\mathcal{A}$, it is clear by the definition that $\operatorname{\mathcal{O}}^{-1}(\mathcal{U})=^{\bot}\mathcal{U}^{c}$. Then [Kr, Corollary 4.3] and section˜4 imply that $\operatorname{\mathcal{O}}^{-1}(\mathcal{U})$ is a localizing subcategory of finite type of $\mathcal{A}$. To prove $\mathcal{X}=\operatorname{\mathcal{O}}^{-1}(\operatorname{\mathcal{O}}(\mathcal{X}))$, clearly $\mathcal{X}\subset\operatorname{\mathcal{O}}^{-1}(\operatorname{\mathcal{O}}(\mathcal{X}))$. For the other side, by the previous argument, $\operatorname{\mathcal{O}}^{-1}(\operatorname{\mathcal{O}}(\mathcal{X}))$ is a localizing subcategory of finite type of $\mathcal{A}$. Thus, for every $M\in\operatorname{\mathcal{O}}^{-1}(\operatorname{\mathcal{O}}(\mathcal{X}))$, we have $M=\underset{\rightarrow}{\rm lim}M_{i}$ where each $M_{i}$ belongs to $\operatorname{\mathcal{O}}^{-1}(\operatorname{\mathcal{O}}(\mathcal{X}))\cap\operatorname{fp-}\mathcal{A}$ by [Kr, Lemma 2.3]. Then $\operatorname{\mathcal{O}}(M_{i})\subset\operatorname{\mathcal{O}}(\mathcal{X})$ for each $i$. Fixing $i$, since $\operatorname{\mathcal{O}}(M_{i})$ is a quasi-compact open subset of $\operatorname{Zg}\mathcal{A}$ by [Kr, Corollary 4.6], there exists $N\in\mathcal{X}\cap\operatorname{fp-}\mathcal{A}$ such that $\operatorname{\mathcal{O}}(M_{i})\subset\operatorname{\mathcal{O}}(N)$ and hence it follows from [H, Corollary 3.12] that $M_{i}\in\surd{N}$, where $\surd{N}$ is the smallest Serre subcategory of $\operatorname{fp-}\mathcal{A}$ containing $N$. Clearly $\surd{N}\subset\mathcal{X}$ and hence $M_{i}\in\mathcal{X}$. Finally, this forces that $M\in\mathcal{X}$ as $M=\underset{\rightarrow}{\rm lim}M_{i}$. It is clear that $\operatorname{\mathcal{O}}(\operatorname{\mathcal{O}}^{-1}(\mathcal{U}))\subset\mathcal{U}$. On the other hand, for every $I\in\mathcal{U}$, there exists a finitely presented object $M$ such that $I\in\mathcal{O}(M)\subset\mathcal{U}$. This implies that $M\in\operatorname{\mathcal{O}}^{-1}(\mathcal{U})$ and so the previous argument forces that $\operatorname{\mathcal{O}}M\subset\operatorname{\mathcal{O}}(\operatorname{\mathcal{O}}^{-1}(\mathcal{U}))$. Therefore $I\in\operatorname{\mathcal{O}}(\operatorname{\mathcal{O}}^{-1}(\mathcal{U}))$; and hence $\mathcal{U}=\operatorname{\mathcal{O}}(\operatorname{\mathcal{O}}^{-1}(\mathcal{U}))$. ∎
5. In the case of commutative coherent rings
In this section, we assume that $A$ is a commutative coherent ring and we denote the category of $A$-modules by $\operatorname{Mod}A$. Also, we denote $\operatorname{ZASpec}\operatorname{Mod}A$, $\operatorname{Zg}\operatorname{Mod}A$ and $\operatorname{fp-}\operatorname{Mod}A$ by $\operatorname{ZASpec}A$, $\operatorname{Zg}A$ and $\operatorname{fp-}A$, respectively.
We recall that a subspace $Y$ of a topological space $X$ is retract provided that there exists a continuous function $r:X\to Y$ such that $r(y)=y$ for all $y$ in $Y$. Let $E$ be any indecomposable injective $A$-module. Set $\mathfrak{p}_{E}$ to be the sum of annihilator ideals of nonzero elements of $E$. As $E$ is uniform, in view of [P, 9.2], it is easy to check that $\mathfrak{p}_{E}$ is a prime ideal of $A$. If we identify every $\alpha\in\operatorname{ZASpec}A$ with $E(\alpha)$ in $\operatorname{Zg}A$, then we have:
Lemma 5.1.
$\operatorname{ZASpec}\mathcal{A}$ is a retract subspace of $\operatorname{Zg}\mathcal{A}$.
Proof.
We define $\theta:\operatorname{Zg}A\rightarrow\operatorname{ZASpec}A$ given by $E\rightarrow E(A/{\mathfrak{p}_{E}})$. If $E(\alpha)\in\operatorname{ZASpec}\mathcal{A}$, then there exists a prime ideal $\mathfrak{p}=\mathfrak{p}_{E(\alpha)}$ of $A$ such that $\alpha=\overline{A/\mathfrak{p}}$ by section˜4. Hence $E(\alpha)=E(A/\mathfrak{p})$ so that $\theta(E(\alpha))=E(\alpha)$. For every finitely presented $A$-module $M$, we show that $\theta^{-1}(\operatorname{ASupp}M)=\mathcal{O}(M)$. If $E\in\theta^{-1}(\operatorname{ASupp}M)$, we have $\theta(E)=E(A/\mathfrak{p}_{E})\in\operatorname{ASupp}M$ so that $E(A/\mathfrak{p}_{E})\in\mathcal{O}(M)$ by section˜3. It now follows from [GP, Theorem 1.4] that $E\in\mathcal{O}(M)$. Conversely if $E\in\mathcal{O}(M)$, then $\theta(E)=E(A/\mathfrak{p}_{E})\in\operatorname{ASupp}M$ by [GP, Theorem 1.4]. Therefore $E\in\theta^{-1}(\operatorname{ASupp}M)$. ∎
Lemma 5.2.
The set $\mathcal{F}\mathcal{I}=\{\operatorname{ASupp}A/\mathfrak{a}|\hskip 2.84544pt\mathfrak{a}$ is a finitely generated ideal of A}A\} forms a basis for the topology $\operatorname{ZASpec}A$.
Proof.
We observe that $\operatorname{ASpec}A=\operatorname{ASupp}A$. For any finitely generated ideals $\mathfrak{a}$ and $\mathfrak{b}$ of $A$ and $\alpha\in\operatorname{ASupp}A/\mathfrak{a}\cap\operatorname{ASupp}A/\mathfrak{b}$. By section˜4, there exists a prime ideal $\mathfrak{p}$ such that $\overline{A/\mathfrak{p}}=\alpha$. It follows from section˜3 that $\operatorname{Hom}_{A}(A/\mathfrak{a},E(A/\mathfrak{p}))\neq 0$ and $\operatorname{Hom}_{A}(A/\mathfrak{b},E(A/\mathfrak{p}))\neq 0$. This implies that $\mathfrak{p}\in V(\mathfrak{a})\cap V(\mathfrak{b})=V(\mathfrak{a}+\mathfrak{b})$; and hence we have $\alpha\in\operatorname{ASupp}A/(\mathfrak{a}+\mathfrak{b})$. Therefore $\operatorname{ASupp}A/\mathfrak{a}\cap\operatorname{ASupp}A/\mathfrak{b}\subseteq\operatorname{ASupp}A/(\mathfrak{a}+\mathfrak{b})$ and an easy argument implies that $\operatorname{ASupp}A/\mathfrak{a}\cap\operatorname{ASupp}A/\mathfrak{b}=\operatorname{ASupp}A/(\mathfrak{a}+\mathfrak{b})$. On the other hand, assume that $M$ is a finitely presented $A$-module. Then $M$ is finitely generated, say $M=<x_{1},\dots,x_{n}>$. Set $M_{0}=0$ and $M_{k}=<x_{1},\dots x_{k}>$ for $1\leq k\leq n$. Then $M_{i}/M_{i-1}$ is cyclic and finitely presented for each $i$. Thus there exist finitely generated ideals $\mathfrak{a}_{i}$ such that $M_{i}/M_{i-1}\cong A/\mathfrak{a}_{i}$ for all $1\leq i\leq n$. It is clear that $\operatorname{ASupp}M=\bigcup_{i=1}^{n}\operatorname{ASupp}A/\mathfrak{a}_{i}$. ∎
For a spectral topological space $X$, Hochster [Ho] endows the underlying set with a new, dual topological by defining its open subsets as those of the form $Y=\bigcup_{i\in\varOmega}Y_{i}$, where $X\setminus Y_{i}$ is a quasi-compact open subset of $X$ for each $i\in\varOmega$. The symbol $X^{\star}$ denotes $X$ with the new topology. We write $\operatorname{Spec}^{\star}A$ for $(\operatorname{Spec}A)^{\star}$.
Proposition 5.3.
The maps
$\operatorname{Spec}^{\star}A\supseteq\mathcal{V}\stackrel{{\scriptstyle\mathcal{O}}}{{\mapsto}}\mathcal{O}_{\mathcal{V}}=\{\alpha=\overline{A/\mathfrak{p}}\in\operatorname{ZASpec}A|\hskip 2.84544pt\mathfrak{p}\in\mathcal{V}\}\subseteq\operatorname{ZASpec}A\hskip 4.26773ptand$
$\operatorname{ZASpec}A\supseteq\hskip 2.84544pt\mathcal{O}\stackrel{{\scriptstyle g}}{{\mapsto}}\mathcal{V}_{\mathcal{O}}=\{\mathfrak{p}\in\operatorname{Spec}A|\hskip 2.84544pt\overline{A/\mathfrak{p}}\in\mathcal{O}\}\subseteq\operatorname{Spec}^{\star}A$
establish a bijective correspondence between the lattices of open subsets of $\operatorname{Spec}^{\star}A$ and those of $\operatorname{ZASpec}A$.
Proof.
For every finitely generated ideal $\mathfrak{a}$ of $A$, it is known that $\operatorname{Spec}A\setminus V(\mathfrak{a})$ is a qausi-compact open subset of $\operatorname{Spec}A$. Therefore for every open subset $\mathcal{V}$ of $\operatorname{Spec}^{\star}A$, we have $\mathcal{V}=\bigcup_{\gamma\in\varGamma}V(\mathfrak{a}_{\gamma})$ where $\mathfrak{a}_{\gamma}$ ranges over finitely generated ideals of $A$. By the same argument used in the proof of section˜5, we have $\mathcal{O}_{V(\mathfrak{a}_{\gamma})}=\operatorname{ASupp}A/\mathfrak{a}_{\gamma}$ for all $\gamma\in\varGamma$; and hence $\{\overline{A/\mathfrak{p}}\in\operatorname{ZASpec}A|\hskip 2.84544pt\mathfrak{p}\in\mathcal{V}\}=\bigcup_{\gamma\in\varGamma}\operatorname{ASupp}A/\mathfrak{a}_{\gamma}$ is an open subset of $\operatorname{ZASpec}A$. Let $\mathcal{O}$ be an open subset of $\operatorname{ZASpec}\mathcal{A}$. By section˜5, we have $\mathcal{O}=\bigcup_{\varLambda}\operatorname{ASupp}A/\mathfrak{a}_{\lambda}$, where $\mathfrak{a}_{\lambda}$ ranges over finitely generated ideals. Hence $\{\mathfrak{p}\in\operatorname{Spec}A|\hskip 2.84544pt\overline{A/\mathfrak{p}}\in\mathcal{O}\}=\bigcup_{\lambda\in\varLambda}V(\mathfrak{a}_{\lambda})$ is a open subset of $\operatorname{Spec}^{\star}A$. Now, it is easy to see that $\mathcal{O}_{\mathcal{V}_{\mathcal{O}}}=\mathcal{O}$ and $\mathcal{V}_{\mathcal{O}_{\mathcal{V}}}=\mathcal{V}$. ∎
Lemma 5.4.
The set $\{\operatorname{\mathcal{O}}(A/\mathfrak{a})|\hskip 2.84544pt\mathfrak{a}$ is a finitely generated ideal of $\mathcal{A}\}$ forms a basis for the topology $\operatorname{Zg}A$.
Proof.
We show that $\operatorname{Zg}A=\bigcup_{\mathfrak{a}\in\text{fg-}A}\mathcal{O}(A/\mathfrak{a})$. Any indecomposable injective module $E$ contains a cyclic submodule $A/\mathfrak{c}$. This implies that $E\in\operatorname{\mathcal{O}}(A/\mathfrak{c})$. If $\mathfrak{c}=0$, there in nothing to prove. Otherwise, $\mathfrak{c}$ contains a finitely generated ideal $\mathfrak{d}$ so that $E\in\mathcal{O}(A/\mathfrak{d})$. The rest of the proof is similar to that in section˜5. ∎
A set $\mathfrak{F}$ of ideals of $A$ is said to be a Gabriel topology on $A$ if it satisfies:
(1). If $\mathfrak{a}\in\mathfrak{F}$ and $\mathfrak{a}\subset\mathfrak{b}$, then $\mathfrak{b}\in\mathfrak{F}$.
(2). If $\mathfrak{a}$, $\mathfrak{b}$ belong to $\mathfrak{F}$, then $\mathfrak{a}\cap\mathfrak{b}\in\mathfrak{F}$.
(3). If $\mathfrak{a}\in\mathfrak{F}$ and $a\in A$, then $(\mathfrak{a}:a)\in\mathfrak{F}$.
(4). If $\mathfrak{a}$ is an ideal of $A$ and there exists $\mathfrak{b}\in\mathfrak{F}$ such that $(\mathfrak{a}:b)\in\mathfrak{F}$ for every $b\in\mathfrak{b}$, then $\mathfrak{a}\in\mathfrak{F}$.
A Gabriel topology $\mathfrak{F}$ is said to have a basis of finitely generated ideals, if any ideal $\mathfrak{a}\in\mathfrak{F}$ contains a finitly generated ideal $\mathfrak{b}\in\mathfrak{F}$. Let $E$ be an indecomposable injective module in $\operatorname{Zg}A$. Clearly, the Gabriel topology
$\mathfrak{F}_{E}=\{\mathfrak{a}|\hskip 2.84544ptE\notin\operatorname{\mathcal{O}}(A/\mathfrak{a})\}$
is associated to $\{M\in\operatorname{Mod}A|\hskip 2.84544pt\operatorname{Hom}_{A}(M,E)=0\}$, the localizing subcategory cogenerated by $E$. More generally, let $\mathcal{E}$ be a subset of $\operatorname{Zg}A$. To $\mathcal{E}$, we associated a Gabriel topology
$\mathfrak{F}_{\mathcal{E}}=\bigcap_{E\in\mathcal{E}}\mathfrak{F}_{E}=\{\mathfrak{a}|\hskip 2.84544pt\operatorname{\mathcal{O}}(A/\mathfrak{a})\cap\mathcal{E}=\emptyset\}.$
On the other hand, for every Gabriel topology $\mathfrak{F}$ on $A$, we associate
$D(\mathfrak{F})=\{E\in\operatorname{Zg}A|\hskip 2.84544ptE\notin\operatorname{\mathcal{O}}(A/\mathfrak{a})\hskip 2.84544pt{\rm for\hskip 2.84544ptevery}\hskip 2.84544pt\mathfrak{a}\in\mathfrak{F}\}=\operatorname{Zg}A\setminus\bigcup_{\mathfrak{a}\in\mathfrak{F}}\operatorname{\mathcal{O}}(A/\mathfrak{a}).$
Given a Gabriel topology $\mathfrak{F}$, an $A$-module $M$ is said to be $\mathfrak{F}$-discrete if $\operatorname{Ann}(x)\in\mathfrak{F}$ for all $x\in M$. The following theorem provides a new classification of Gabriel topologies on $A$.
Lemma 5.5.
Let $D$ be a closed subset of $\operatorname{Zg}A$. Then $\mathfrak{F}_{D}$ has a basis of finitely generated ideals of $A$.
Proof.
We observe that $\operatorname{Zg}A\setminus D$ is an open subset of $\operatorname{Zg}A$; and hence by section˜5, there exists a set of finitely generated ideals $\{\mathfrak{b}_{\lambda}|\hskip 2.84544pt\lambda\in\varLambda\}\subseteq\mathfrak{F}_{D}$ such that $\operatorname{Zg}A\setminus D=\bigcup_{\lambda\in\varLambda}\mathcal{O}(A/\mathfrak{b}_{\lambda})$. Now, given an ideal $\mathfrak{a}\in\mathfrak{F}_{D}$, by the definition, $\mathcal{O}(A/\mathfrak{a})\subseteq\bigcup_{\lambda\in\varLambda}\mathcal{O}(A/\mathfrak{b}_{\lambda})$ so that $A/\mathfrak{a}$ is a quotient of $A/\mathfrak{c}\in\sqrt{\langle A/\mathfrak{b}_{\lambda}|\hskip 2.84544pt\lambda\in\varLambda\rangle}$ by section˜4. This implies that $\mathfrak{c}\subseteq\mathfrak{a}$ and there exists a finite filtration of submodules of $A/\mathfrak{c}$
$0=X_{0}\subset X_{1}\subset\dots\subset X_{n-1}\subset X_{n}=A/\mathfrak{c}$
such that each $X_{i}/X_{i-1}$ is a subquotient of $A/\mathfrak{b}_{\lambda_{i}}$ for some $\lambda_{i}\in\varLambda$. Setting $\mathfrak{b}=\mathfrak{b}_{\lambda_{1}}\mathfrak{b}_{\lambda_{2}}\dots\mathfrak{b}_{\lambda_{n}}$, it follows that $\mathfrak{b}\subset\mathfrak{c}$ and $\mathfrak{b}\in\mathfrak{F}_{D}$ by [St, Chap VI, p. 147, Lemma 5.3]. ∎
There are many classifications of Gabriel topologies on $A$ (e.g., see [G, St]). For commutative coherent rings, the following theorem provides a new classification of Gabriel topologies on $A$. The set of finitely generated modules is denoted by fg-$A$.
Theorem 5.6.
There are bijective correspondence between:
$\rm(1)$} the lattice of open subsets of $\operatorname{ZASpec}A$,
$\rm(2)$} the lattice of closed subsets of $\operatorname{Zg}A$,
$\rm(3)$} the lattice of Gabriel topologies having bases of finitely generated ideals of $A$,
$\rm(4)$} The lattice of localizing subcategory of finite type of $\operatorname{Mod}A$.
These bijections are defined as follows:
${\rm(1)}\longleftrightarrow{\rm(2)}$: $\operatorname{ASupp}A/\mathfrak{a}\longleftrightarrow\operatorname{Zg}A\setminus\mathcal{O}(A/\mathfrak{a})$ for any finitely generated ideal $\mathfrak{a}$.
${\rm(2)}\longleftrightarrow{\rm(3)}$: $D\stackrel{{\scriptstyle\mathfrak{F}}}{{\mapsto}}\mathfrak{F}_{D}$ and $\mathfrak{F}\stackrel{{\scriptstyle D}}{{\mapsto}}D(\mathfrak{F})$,
${\rm(3)}\longleftrightarrow{\rm(4)}$: $\mathfrak{F}\mapsto\mathcal{X}_{\mathfrak{F}}=\overrightarrow{\sqrt{\langle A/\mathfrak{a}|\hskip 2.84544pt\mathfrak{a}\in\mathfrak{F}\cap\operatorname{fg-}A\rangle}}$ and $\mathcal{X}\mapsto\mathfrak{F}_{\mathcal{X}}=\{\mathfrak{a}|\hskip 2.84544ptA/\mathfrak{a}\in\mathcal{X}\}$.
Proof.
${\rm(1)}\longleftrightarrow{\rm(2)}$ follows from section˜3, section˜4, section˜5 and section˜5.
${\rm(2)}\longleftrightarrow{\rm(3)}$. Let $\mathfrak{F}$ be a Gabriel topology on $A$ having a basis of finitely generated ideal of $A$ and $D$ be a closed subset of $\operatorname{Zg}A$. Then section˜5 implies that $\mathfrak{F}_{D}$ has a basis of finitely generated ideals and $D(\mathfrak{F})=\operatorname{Zg}A\setminus\bigcup_{\mathfrak{a}\in\mathfrak{F}}\operatorname{\mathcal{O}}(A/\mathfrak{a})=\operatorname{Zg}A\setminus\bigcup_{\mathfrak{a}\in\mathfrak{F}\cap\operatorname{fg-}A}\operatorname{\mathcal{O}}(A/\mathfrak{a})$ is a closed subset of $\operatorname{Zg}A$ by section˜5. We show that $\mathfrak{F}_{D(\mathfrak{F})}=\mathfrak{F}$. Given $\mathfrak{a}\in\mathfrak{F}$, the definition implies that $\operatorname{Hom}_{A}(A/\mathfrak{a},D(\mathfrak{F}))=0$ so that $\mathfrak{a}\in\mathfrak{F}_{D(\mathfrak{F})}$. Conversely, if $\mathfrak{b}\in\mathfrak{F}_{D(\mathfrak{F})}$, then $\operatorname{Hom}_{A}(A/\mathfrak{b},D(\mathfrak{F}))=0$ so that $\operatorname{\mathcal{O}}(A/\mathfrak{b})\subseteq\bigcup_{\mathfrak{a}\in\mathfrak{F}}\operatorname{\mathcal{O}}(A/\mathfrak{a})$. It then follows from section˜5 that $\operatorname{\mathcal{O}}(A/\mathfrak{b})\subseteq\bigcup_{\mathfrak{a}\in\mathfrak{F}\cap{\text{fg-}}A}\operatorname{\mathcal{O}}(A/\mathfrak{a})=\mathcal{O}(\overrightarrow{\sqrt{\langle A/\mathfrak{a}|\hskip 2.84544pt\mathfrak{a}\in\mathfrak{F}\cap{\text{fg-}}A\rangle}})$. Now, section˜4 implies that $A/\mathfrak{b}\in\overrightarrow{\sqrt{\langle A/\mathfrak{a}|\hskip 2.84544pt\mathfrak{a}\in\mathfrak{F}\cap{\text{fg-}}A\rangle}}$. Therefore, $A/\mathfrak{b}$ is a quotient of some $A/\mathfrak{a}$ for $\mathfrak{a}\in\mathfrak{F}\cap{\text{fg-}}A$. This implies that $\mathfrak{a}\subseteq\mathfrak{b}$ so that $\mathfrak{b}\in\mathfrak{F}$. To prove $D(\mathfrak{F}_{D})=D$, since $\operatorname{Zg}A\setminus D$ is an open subset of $\operatorname{Zg}\mathcal{A}$, by section˜5, there exists a set of finitely generated ideals $\{\mathfrak{a}_{\lambda}|\hskip 2.84544pt\lambda\in\varLambda\}$ of $A$ such that $\operatorname{Zg}A\setminus D=\bigcup_{\lambda\in\varLambda}\operatorname{\mathcal{O}}(A/\mathfrak{a}_{\lambda})$. The inclusion $D\subseteq D(\mathfrak{F}_{D})$ is clear. Conversely, for any $E\in D(\mathfrak{F}_{D})$, since $\mathfrak{a}_{\lambda}\in\mathfrak{F}_{D}$ for any $\lambda\in\varLambda$, we have $E\notin\operatorname{\mathcal{O}}(A/\mathfrak{a}_{\lambda})$ for any $\lambda\in\varLambda$. Consequently $E\in D$.
${\rm(3)}\longleftrightarrow{\rm(4)}$. Given a Gabriel topology $\mathfrak{F}$ with a basis of finitely generated ideals of $A$, it follows from [Kr, Theorem 2.8] that $\mathcal{X}_{\mathfrak{F}}$ is a localizing subcategory of finite type of $\mathcal{A}$. Moreover, for a localizing subcategory $\mathcal{X}$ of $\operatorname{Mod}A$ of finite type, it follows from [St, Chap VI, p. 146, Theorem 5.1] that $\mathfrak{F}_{\mathcal{X}}$ is a Gabriel topology on $A$. If $\mathfrak{a}\in\mathfrak{F}_{\mathcal{X}}$, then we have $A/\mathfrak{a}=\underset{\rightarrow}{\rm lim}M_{i}$ is the direct limit of objects $M_{i}$ in $\mathcal{X}\cap\operatorname{fp-}A$. Then $\mathfrak{a}$ contains a finitely generated ideal $\mathfrak{b}$ such that $A/\mathfrak{b}\in\mathcal{X}$. Hence $\mathfrak{F}_{\mathcal{X}}$ has a basis of finitely generated ideals of $A$. We first show that $\mathfrak{F}_{\mathcal{X}_{\mathfrak{F}}}=\mathfrak{F}$. Given $\mathfrak{a}\in\mathfrak{F}$, it contains a finitely generated ideal $\mathfrak{b}\in\mathfrak{F}$ so that $A/\mathfrak{b}\in\mathcal{X}$. Then $A/\mathfrak{a}\in\mathcal{X}_{\mathfrak{F}}$; and hence $\mathfrak{a}\in\mathfrak{F}_{\mathcal{X}_{\mathfrak{F}}}$. Conversely, if $\mathfrak{a}\in\mathfrak{F}_{\mathcal{X}_{\mathfrak{F}}}$, then $A/\mathfrak{a}\in\mathcal{X}_{\mathfrak{F}}$. By a similar argument as mentioned above, $\mathfrak{a}$ contains a finitely generated ideal $\mathfrak{b}\in\mathfrak{F}$ so that $\mathfrak{a}\in\mathfrak{F}$. Now, we show that $\mathcal{X}_{\mathfrak{F}_{\mathcal{X}}}=\mathcal{X}$. Given $M\in\mathcal{X}$, since $\mathcal{X}$ and $\mathcal{X}_{\mathfrak{F}_{\mathcal{X}}}$ are localizing subcategory of finite type of $\mathcal{A}$, we may assume that $M$ is finitely presented. Assume that $M=<x_{1},\dots,x_{n}>$, $M_{0}=0$ and $M_{k}=<x_{1},\dots x_{k}>$ for $1\leq k\leq n$. Then $M_{i}/M_{i-1}=A/\mathfrak{a}_{i}\in\mathcal{X}$ for $1\leq i\leq n$. Hence $M_{i}/M_{i-1}\in\mathcal{X}_{\mathfrak{F}_{\mathcal{X}}}$ for $1\leq i\leq n$; and consequently $M\in\mathcal{X}_{\mathfrak{F}_{\mathcal{X}}}$. ∎